1- A power line carries a current of 81A west along the tops of 8.5-m-high poles.

a) How does this compare with the Earth's magnetic field of about 12 G?

b) here would the wire's magnetic field cancel the Earth's field?
Answer #1

1- A power line carries a current of 81A west along the tops of 8.5-m-high poles.

a) How does this compare with the Earth's magnetic field of about 12 G?

b) here would the wire's magnetic field cancel the Earth's field?this question has nothing to do with
magnetic force F = I x (length) x B x sin90, another formula should be used.
The magnetic field produced by a long wire
B = (?o/4?)(2I/r) = 10^(-7)*2*81/8.5 T = 2.2 x 10^-6 T = 0.022 G
B / B(earth) = 0.044